# Neural Approaches for Multi-Objective Routing on Multigraphs

### v2

@@ 4.3 Training Algorithm
Input: Preference distribution $\Lambda$, instance distribution $S$, batch size $B$, number of samples per instance head 1, $K_{1}$, and head 2, $K_{2}$
@@
6:  Calculate gradients $\nabla J_{1}(\tilde{\theta})$, $\nabla J_{2}(\theta)$ with (LABEL:eq:reinforce)
7:  $\tilde{\theta}\leftarrow$ADAM$(\tilde{\theta},\nabla J_{1}(\tilde{\theta}))$, $\theta\leftarrow$ ADAM$(\theta,\nabla J_{2}(\theta))$,
In addition, we utilize a form of curriculum learning to speed up training for both models. We start by training on small graph sizes and then gradually increase the problem size. For GMS-DH in the multigraph setting, we also start by training the routing-head on simple graphs to ensure that the approximation (18) is accurate even during the initial epochs of multigraph training. Further details regarding the curriculum learning can be found in Appendix C.
5 Experimental Studies
5.1 Benchmarking Setup
Problems
In the asymmetric simple graph case we consider two standard routing problems: The Multi-Objective Traveling Salesman Problem (MOTSP) and Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP). Specifically, we consider bi-objective variants of these. In the MOTSP, we sample two distances for each node pair. These are summarized over the tour to yield the objective values. Regarding the distances, we consider both the TMAT (Kwon et al. 2021) and XASY distributions (Gaile et al. 2022), where the former obeys the triangle inequality and the latter does not - making it more difficult. For the MOCVRP, we sample one distance according to these distributions. We then let the first objective be given by the total distance of a tour and the second objective by the makespan, which is the longest trip for a single vehicle. This is a standard benchmark problem in literature (Lin, Yang, and Zhang 2022).
@@ 5.1 Benchmarking Setup
In Appendix D, we provide more in-depth explanations of the problems and distributions. We also provide results on Euclidean problems in Appendix G.
@@
Table 1: MOTSP, MOCVRP and MGMOTSP with different instance sizes (50, 100) and distributions. Runtime is the total time for solving 200 instances. The best results (excluding LKH, our baseline) are in bold and the second best are underlined.
Baselines
For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun 2017) and Google OR-tools (Perron and Furnon 2019). We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi and Murata 1998), MOEA/D (Zhang and Li 2007), NSGA-II (Deb et al. 2002) and NSGA-III (Blank, Deb, and Roy 2019). Below, we only show the results for the MOEA that performs best in each scenario.
Finally, we also design a neural baseline based on the MatNet-architecture (Kwon et al. 2021) as encoder and a preference-conditioned MP network as decoder. We call this MatNet-Based Model (MBM) and our motivation is that it is a method that can be used almost ”off-the-shelf” by combining components from literature. In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization following Proposition 1.
In Appendix B and E, we provide further details about the neural benchmark and MOEAs respectively. In Appendix G, we also show results for all MOEAs as well as some other weaker heuristics.
Model Settings
We train all models for $200$ epochs using 100 000100000100\,000100 000 randomly generated instances per epoch with the ADAM optimizer (Kingma and Ba 2015). We use the batch size $B=64$ and the learning-rate $\eta=10^{-4}$. For GMS-EB we let the number of layers be $L=6$, while GMS-DH has $L_{1}=5$ and $L_{2}=2$. For both models we set the embedding dimension to 128 and use $8$ attention heads in the GREAT layers and MP decoders. Regarding the number of samples in the decoding, we set $K_{1}=4$ for the selection-decoder during training and $K_{1}=1$ during inference. In the routing-decoder, we always set $K_{2}$ to the problem size as customary in the POMO framework (Kwon et al. 2020). Finally, we augment the models during inference with a factor of $8$. For MBM, this is done using the original MatNet-augmentation (Kwon et al. 2021), while for GMS we use the scaling-augmentation of Lischka et al. (2025).
Metrics
We utilize the Hypervolume (HV) metric (Zitzler et al. 2003) to evaluate the performance of the methods. Formally, for a given set of points $S=\{y_{1},\dots,y_{n}\}\subset\mathbb{R}^{m}$ and a reference point $r\in\mathbb{R}^{m}$ that is dominated by all points in $S$, the HV is defined as:
$\text{HV}(S,r)=\Lambda\left(\bigcup_{i=1}^{n}[y_{i},r]\right),$ (20)
where $\Lambda$ denotes the Lebesgue measure in $\mathbb{R}^{m}$, and $[y_{i},\mathbf{r}]$ represents the axis-aligned hyperrectangle bounded by $y_{i}$ and $r$. Intuitively, a higher HV generally indicates better performance, since the points in the obtained pareto front are further from the reference.
We report the normalized HV, which is scaled to lie in the interval $[0,1]$, averaged over 200 test instances. For the neural models we use 101 preferences during inference, which is also the number of subproblems for the benchmarks relying on scalarization. Apart from the HV, we provide the HV gap compared to the best-performing method as well as the inference time. In Appendix C, we also report some training times for the learning-based methods.
Hardware
We conducted training and inference for the learning-based methods using a single NVIDIA Tesla A40 GPU with 48GB of VRAM. For the non-learning-based methods, evaluations were performed on an Intel Xeon Gold 6338 CPU and, to enable a more fair comparison, we report execution time based on solving multiple instances in parallel. Akin to Kwon et al. (2021), we set the number of parallel processes to 8.
5.2 Performance on Benchmark Problems
@@
The poor performance of GMS-DH PP in Table 1 highlights the importance of explicitly designing for the multigraph structure, even when a theoretically sound pruning strategy is available. Since GMS-DH PP requires re-running the encoder for each preference, it is approximately 3$\times$ slower than GMS-DH. We also hypothesize that the pre-pruning step shifts the data distribution in a way that GMS-DH is ill-equipped to model, resulting in performance that is worse than MBM with the same pruning strategy. Indeed, in our experience GMS-DH PP tends to produce Pareto fronts with poor diversity, clustered around central preference regions
@@ 5.3 Zero-Shot Generalization
5.4 More Difficult Multigraph Problem
Besides MBM with linear pre-pruning and MOEAs, we benchmark against a simplified version of GMS-DH in order to ablate its ability to learn which edges are optimal. In this simplified version, we replace the selection-decoder with a simple function that removes edges with suboptimal linear cost. Compared to pre-pruning the multigraph before the encoder, this approach is smarter, as it removes the need to reencode. Furthermore, this simple variant actually performs quite well on the MGMOTSP, which we show and discuss in Appendix F.
@@ 5.4 More Difficult Multigraph Problem
Table 3: MGMOTSPTW on the FLEX2 and FIX2 distributions with 50 nodes. Runtime is for solving 200 instances. In GMS-DH Simple, the selection-decoder is replaced with a simple pruning function.
@@ Acknowledgements
Appendix A Extended Related Work
@@
In this section, we complement the literature review in the main text with a more thorough discussion of related work.
A.1 Single-Objective Learning for Routing
@@
In the autoregressive paradigm, Vinyals, Fortunato, and Jaitly (2015) were early to propose learning solutions to the TSP. They utilized an oracle solver for supervised learning, and Bello et al. (2017) later introduced reinforcement learning for training. By using attention models and inherent symmetries in the problem, first Kool, van Hoof, and Welling (2018) and then Kwon et al. (2020) obtained greatly improved optimality gaps on various routing problems. Regarding asymmetric problems, Kwon et al. (2021) proposed MatNet, a mixed-attention architecture working on bi-partite graphs. More recent works are Jin et al. (2023); Drakulic et al. (2023); Chen et al. (2024); Bi et al. (2024). These papers have improved results, can handle constraints better and can scale to even larger problems than previously. Another recent notable contribution was made by Drakulic, Michel, and Andreoli (2025), who proposed a jointly trained generalist model to solve a multitude of combinatorial optimization problems, including examples from routing.
@@
In the non-autoregressive category, some works are Joshi, Laurent, and Bresson (2019); Fu, Qiu, and Zha (2021); Qiu, Sun, and Yang (2022); Sun and Yang (2023). Typically, these methods require more sophisticated procedures than simple sampling in the decoding, such as Monte–Carlo tree search, to perform competitively, but scale better to large instances due to lower space and time complexity. Thus, an interesting recent work combining autoregressive and non-autoregressive construction to solve large instances (up to $10^{5}$ nodes for the TSP) while maintaining performance is Ye et al. (2024).
@@
Besides constructive approaches, improvement-based methods usually obtain great results, but with longer running-times, by refining initial solutions. Some examples are Chen and Tian (2019); Lu, Zhang, and Yang (2020); Hottung and Tierney (2020); Hudson et al. (2022).
A.2 Multi-Objective Learning for Routing
@@
Early works treating MO routing using learning utilize multiple models, where one model is responsible for one preference between the objectives (Li, Zhang, and Wang 2021; Wu, Wang, and Zhang 2020). This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott 2005). Pareto-set learning for multi-objective routing using a single model was proposed by Lin, Yang, and Zhang (2022). They used the hyper-networks of Navon et al. (2021) to condition the decoder on the preference. Recent studies that feature improved results with a similar conditioning method include Fan et al. (2025) and Lu et al. (2024). An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems. Two representative works in this direction are Zhang et al. (2021) and Chen et al. (2023a).
@@
A.3 Routing on Multigraphs
Successful approaches to tackle MO shortest path problems on multigraphs include genetic algorithms (Beke, Weiszer, and Chen 2021), A-related methods (Weiszer, Burke, and Chen 2020) and hybrid approaches combining A with learning (Liu, Chen, and Weiszer 2022). These methods all aim to find a Pareto set of shortest paths between two destinations.
For combinatorial routing problems such as VRPs on multigraphs, heuristics and exact solvers have been proposed. The first study in this regard was conducted by Garaix et al. (2010), who designed a simple heuristic and branch-and-price method for a single-objective on-demand transportation problem. Other works for single-objective constrained problems or multi-objective problems are Lai, Caliskan Demirag, and Leung (2016); Ben Ticha et al. (2019); Tikani, Setak, and Demir (2021). In addition, Ben Ticha et al. (2017) clearly concluded that a multigraph setup yields significantly lower costs compared to a simple graph representation for the capacitated VRP with time-windows. At the same time, the computation time increases significantly under this perspective, and the authors highlight the need for new heuristics.
Appendix B Architecture Details
@@
B.1 Edge-Based Decoder in GMS-EB
@@
$u_{l}=\begin{cases}c\tanh{\alpha_{l}}&\text{ if }l\text{ connects current and unvisited node}\\ -\infty&\end{cases}$ (24)
@@
B.2 Selection-Decoder in GMS-DH
The task of this decoder is to prune parallel edges in the intermediate stage of GMS-DH, leaving only one between each node pair. Initially, we apply $L_{3}$ GREAT NB layers on the edge embeddings from the encoder to obtain embeddings specialized for this task. In the experiments we set $L_{3}=2$. These layers are not preference conditioned and hence they do not need to be rerun for each preference. Next, given two nodes, we want to compute the probability of retaining each (directed) edge in between. For this purpose, we essentially utilize a multi-pointer score calculation similar to (23). Given a node pair $(i,j)$, and the set of edges in between, $E_{ij}$, let the query be the average embedding according to
@@ B.2 Selection-Decoder in GMS-DH
B.3 MatNet-Based Benchmark Model
We visualize this architecture in Figure 3. The encoder consists of $L$ MatNet-layers (Kwon et al. 2021), while the decoder consists of an MP decoder (Jin et al. 2023) preference-conditioned with a hyper-network.
The MatNet-architecture operates on bipartite graphs. It utilizes mixed-score attention layers together with fully-connected layers and normalization as in standard transformer layers (Vaswani et al. 2017). These attention layers ”mix” internally computed attention scores with external information from e.g. a distance matrix $D$. To accommodate the use of several distance matrices, $D_{1},\dots,D_{f}$, in our multi-objective setting, we follow the idea in Appendix A of the original paper. That is, we mix the internal attention score with all distance matrices using an MLP with $f+1$ inputs.

### v4

@@ 5 Experimental Studies
Input: Preference distribution $\Lambda$, instance distribution $S$,
batch size $B$, number of samples per instance head 1, $K_{1}$,
and head 2, $K_{2}$
@@
[…]
Problems. In the asymmetric simple graph case we consider two standard routing problems: The Multi-Objective Traveling Salesman Problem (MOTSP) and Multi-Objective Capacitated Vehicle Routing Problem (MOCVRP). Specifically, we consider bi-objective variants of these. In the MOTSP, we sample two distances for each node pair. These are summarized over the tour to yield the objective values. Regarding the distances, we consider both the TMAT (Kwon et al., 2021) and XASY distributions (Gaile et al., 2022), where the former obeys the triangle inequality and the latter does not - making it more difficult. For the MOCVRP, we sample one distance according to these distributions. We then let the first objective be given by the total distance of a tour and the second objective by the makespan, which is the longest trip for a single vehicle. The Euclidean version of this problem has been used extensively for benchmarking in previous works (Lin et al., 2022).
In the multigraph setting, we consider the Multigraph MOTSP (MGMOTSP) with two objectives, where we sample two-dimensional edge distances using distributions called FIXxx and FLEXxx. Here, $x$ denotes the number of edges between each node pair. In the FLEX-distribution, we sample edge distances independently and remove edges which are dominated. In the FIX distribution we first sample edges independently, but then rearrange by sorting, so that the an edge that is best in one objective is the worst in the other objective. Thus, in the FIXxx-distribution there are always exactly $x$ edges between each node pair, whereas FLEXxx has at most $x$ edges.
@@
In Appendix E, we provide more in-depth explanations of the problems and distributions. We also provide results on Euclidean problems and tri-objective problems in Appendix H.
Baselines. For the MOTSP and MGMOTSP, we utilize linear scalarization together with strong single-objective methods in the form of LKH3 (Helsgaun, 2017) and Google OR-tools (Perron & Furnon, 2019). We also benchmark against four Multi-Objective Evolutionary Algorithms (MOEAs): MOGLS (Ishibuchi & Murata, 1998), MOEA/D (Zhang & Li, 2007), NSGA-II (Deb et al., 2002) and NSGA-III (Blank et al., 2019). Below, we only show the results for the MOEA that performs best in each scenario.
Finally, we also design a neural baseline based on the MatNet-architecture (Kwon et al., 2021) as encoder and a preference-conditioned MP network as decoder. We call this MatNet-Based Model (MBM) and our motivation is that it is a method that can be used almost ”off-the-shelf” by combining components from literature. In the multigraph setting, this architecture requires us to pre-process the input by pruning edges until we obtain a simple graph, which we do using linear scalarization following Proposition 1.
In Appendix C and F, we provide further details about the neural benchmark and MOEAs respectively. In Appendix H, we also show results for all MOEAs as well as some other weaker heuristics.
Model Settings. We train all models for $200$ epochs using 100 000100\,000 randomly generated instances per epoch with the ADAM optimizer (Kingma & Ba, 2015). We use the batch size $B=64$ and the learning-rate $\eta=10^{-4}$. For each batch, we sample a preference $\lambda=(\lambda_{1},\lambda_{2})$ using $\lambda_{1}\sim\text{Unif}[0,1]$ and $\lambda_{2}=1-\lambda_{1}$. For GMS-EB we let the number of layers be $L=6$, while GMS-DH has $L_{1}=5$ and $L_{2}=2$. Both models have embedding dimension 128 and use $8$ attention heads in the GREAT layers and MP decoders. Regarding the number of samples in the decoding, we set $K_{1}=4$ for the selection-decoder during training and $K_{1}=1$ during inference. In the routing-decoder, we always set $K_{2}$ to the problem size as customary in the POMO framework (Kwon et al., 2020). Finally, we augment the models during inference with a factor of $8$. For MBM, this is done using the original MatNet-augmentation (Kwon et al., 2021), while for GMS we use the scaling-augmentation of Lischka et al. (2025).
Metrics. We utilize the Hypervolume (HV) metric (Zitzler et al., 2003) to evaluate the performance of the methods. Formally, for a given set of points $S=\{y_{1},\dots,y_{n}\}\subset\mathbb{R}^{m}$ and a reference point $r\in\mathbb{R}^{m}$ that is dominated by all points in $S$, the HV is defined as:
$\text{HV}(S,r)=\Lambda\left(\bigcup_{i=1}^{n}[y_{i},r]\right),$ (18)
where $\Lambda$ denotes the Lebesgue measure in $\mathbb{R}^{m}$, and $y_{i},\mathbf{r}]$ represents the axis-aligned hyperrectangle bounded by $y_{i}$ and $r$. Intuitively, a higher HV generally indicates better performance, since the points in the obtained Pareto front are further from the reference.
We report the normalized HV, which is scaled to lie in the interval $[0,1]$, averaged over 200 test instances. For the neural models and benchmarks relying on scalarization, we use 101 preferences linearly spaced between $\lambda=(1,0)$ and $(0,1)$ during inference to obtain the Pareto front. Apart from the HV, we provide the HV gap compared to the best-performing method as well as the inference time. In Appendix D, we also report some training times for the learning-based methods.
Hardware. We conducted training and inference for the learning-based methods using a single NVIDIA Tesla A40 GPU with 48GB of VRAM. For the non-learning-based methods, evaluations were performed on an Intel Xeon Gold 6338 CPU and, to enable a more fair comparison, we report execution time based on solving multiple instances in parallel. Akin to Kwon et al. (2021), we set the number of parallel processes to 8.
5.1 Performance on Benchmark Problems
Table 1: MOTSP, MOCVRP and MGMOTSP with different instance sizes (50, 100) and distributions. Runtime is the total time for solving 200 instances. The best results (excluding LKH, our baseline) are in bold and the second best are underlined. Remark that most neural MO solvers benchmark on instances of roughly this size (Chen et al., 2023b; Fan et al., 2025).
@@ 5.1 Performance on Benchmark Problems
[…]
@@ 5.2 Zero-Shot Generalization
Table 3: MGMOTSPTW on the FLEX2 and FIX2 distributions with 50 nodes. Performance over 200 instances. In GMS-DH Simple, the selection-decoder is replaced with a simple pruning function.
@@
5.3 A Difficult Multigraph Problem
Besides MBM with linear pre-pruning and MOEAs, we benchmark against a simplified version of GMS-DH in order to ablate its ability to learn which edges are optimal. In this simplified version, we replace the selection-decoder with a simple function that removes edges with suboptimal linear cost. Compared to pre-pruning the multigraph before the encoder, this approach is smarter, as it removes the need to reencode. Furthermore, this simple variant actually performs quite well on the MGMOTSP, which we show and discuss in Appendix G.
@@ Acknowledgements
Appendix A Extended Motivation
In this section, we further motivate our approach by explaining why simple transformations of the underlying multigraph, combined with vanilla node-based construction, are unsuitable in the context of vehicle routing problems.
Firstly, it is possible to translate a multigraph into a simple graph by using its line graph, which is shown on the left side of Figure 3. This entails replacing each edge with a node and connecting those new nodes which share endpoints. In our context of fully connected graphs with $M$ parallel edges, this results in $\mathcal{O}(MN^{2})$ nodes in the line graph and $\mathcal{O}(M^{2}N^{3})$ edges. By utilizing node-based construction (that is, selecting one node at a time sequentially) on the new graph with $N$ POMO samples and in $N$ steps, the complexity in the decoding would thus be $\mathcal{O}(MN^{4})$.
Another, related, approach is to introduce virtual end-point nodes for each parallel edge incident to a specific node, which is shown on the right side of Figure 3. Again, this leads to a higher node count, as each edge in the original graph is represented using one virtual node, yielding $\mathcal{O}(MN^{2})$ total nodes. While the complexity of the total number of edges is lower in this case ($\mathcal{O}(MN^{2})$, since we introduce only one edge per virtual node), for standard node-based decoding the complexity remains $\mathcal{O}(MN^{4})$.
One can note that both these transformations combined with node-based construction resemble GMS-EB in terms of complexity and underlying mechanism. At its core, GMS-EB treats each edge as a node in the decoding. However, as GMS-EB works on the original graph, we believe it to be significantly more interpretable. Moreover, GMS-DH obtains lower complexity ($\mathcal{O}(N^{3})$) in its decoding, since it is node-based in the original graph.
Transformation 1Transformation 2Original nodesVirtual nodes
Figure 3: Two possible graph transformations to convert multigraphs into simple graphs. For simplicity, we show both transformations on a graph that is not fully connected.
In summary, while it is possible to work on transformed simple graphs instead of underlying multigraphs, we obtain significant advantages by not doing so. Optimizing architectures on transformed graphs constitutes another research direction that would be interesting to pursue, but is beyond the scope of this work. As shown by the occasionally poor performance of MBM on asymmetric simple graphs in Table 1, it seems unlikely that such an ”off-the-shelf” architecture would work well on much larger graphs induced by transforming multigraphs.
Appendix B Extended Related Work
@@
Here, we complement the literature review in the main text with a more thorough discussion of related work.
B.1 Single-Objective Learning for Routing
@@
In the autoregressive paradigm, Vinyals et al. (2015) were early to propose learning solutions to the TSP. They utilized an oracle solver for supervised learning, and Bello et al. (2017) later introduced reinforcement learning for training. By using attention models and inherent symmetries in the problem, first Kool et al. (2018) and then Kwon et al. (2020) obtained greatly improved optimality gaps on various routing problems. Regarding asymmetric problems, Kwon et al. (2021) proposed MatNet, a mixed-attention architecture working on bi-partite graphs. More recent works are Jin et al. (2023); Drakulic et al. (2023); Chen et al. (2024); Bi et al. (2024). These papers have improved results, can handle constraints better and can scale to even larger problems than previously. Another recent notable contribution was made by Drakulic et al. (2025), who proposed a jointly trained generalist model to solve a multitude of combinatorial optimization problems, including examples from routing.
@@
In the non-autoregressive category, some works are Joshi et al. (2019); Fu et al. (2021); Qiu et al. (2022); Sun & Yang (2023). Typically, these methods require more sophisticated procedures than simple sampling in the decoding, such as Monte–Carlo tree search, to perform competitively, but scale better to large instances due to lower space and time complexity. Thus, an interesting recent work combining autoregressive and non-autoregressive construction to solve large instances (up to $10^{5}$ nodes for the TSP) while maintaining performance is Ye et al. (2024).
@@
Besides constructive approaches, improvement-based methods usually obtain great results, but with longer running-times, by refining initial solutions. Some examples are Chen & Tian (2019); Lu et al. (2020); Hottung & Tierney (2020); Hudson et al. (2022).
B.2 Multi-Objective Learning for Routing
@@
Early works treating MO routing using learning utilize multiple models, where one model is responsible for one preference between the objectives (Li et al., 2021; Wu et al., 2020). This quickly becomes infeasible as the number of models required grows exponentially with the number of objectives (Ehrgott, 2005). Pareto-set learning for multi-objective routing using a single model was proposed by Lin et al. (2022). They used the hyper-networks of Navon et al. (2021) to condition the decoder on the preference. Recent studies that feature improved results with a similar conditioning method include Fan et al. (2025) and Lu et al. (2024). An alternative approach is to use meta-learning to train a single model that can be quickly fine-tuned over a few steps to solve individual subproblems. Two representative works in this direction are Zhang et al. (2021) and Chen et al. (2023a).
@@
B.3 Routing on Multigraphs
We note that apart from vehicle routing problems, which are the focus in this article, much research has been dedicated to MO shortest path problems on multigraphs. Successful approaches for these problems include genetic algorithms (Beke et al., 2021), A-related methods (Weiszer et al., 2020) and hybrid approaches combining A with learning (Liu et al., 2022). These methods all aim to find a Pareto set of shortest paths between two destinations.
Appendix C Architecture Details
@@
C.1 Edge-Based Decoder in GMS-EB
@@
[…]
@@
C.2 Selection-Decoder in GMS-DH
The task of this decoder is to prune parallel edges in the intermediate stage of GMS-DH, leaving only one between each node pair. Initially, we apply $L_{3}$ GREAT NB layers on the edge embeddings from the encoder to obtain embeddings specialized for this task. In the experiments we set $L_{3}=2$. These layers are not preference conditioned and hence they do not need to be rerun for each preference. Next, given two nodes, we want to compute the probability of retaining each (directed) edge in between. For this purpose, we essentially utilize a multi-pointer score calculation similar to equation 21. Given a node pair $(i,j)$, and the set of edges in between, $E_{ij}$, let the query be the average embedding according to
@@ C.2 Selection-Decoder in GMS-DH
C.3 MatNet-Based Benchmark Model
We visualize this architecture in Figure 4. The encoder consists of $L$ MatNet-layers (Kwon et al., 2021), while the decoder consists of an MP decoder (Jin et al., 2023) preference-conditioned with a hyper-network.
The MatNet-architecture operates on bipartite graphs. It utilizes mixed-score attention layers together with fully-connected layers and normalization as in standard transformer layers (Vaswani et al., 2017). These attention layers ”mix” internally computed attention scores with external information from e.g., a distance matrix $D$. To accommodate the use of many edge features represented by distance matrices, $D_{1},\dots,D_{f}$, in our multi-objective setting, we follow the idea in Appendix A of the original paper. That is, we mix the internal attention score with all distance matrices using an MLP with $f+1$ inputs.